the area of a triangle is 27sqft. It's height is 3 times the length of it's base. Find the height and base of the triangle.
27=1/2b(3B)
Just not sure how to solve from here

Answer by rfer(16322)

27=.5b(3b)
27=1.5b^2
27/1.5=b^2
18=b^2
b=4.24
h=3b=12.72